OpenStudy (anonymous):

find the value of x and the value of y for which (x, 14) and (18, y) are points on the graph. y = 0.5x+12 if you ccould just show me how to do this I might get to sleep today.

OpenStudy (anonymous):

i no i don no how 2 do this ether but nice pic

OpenStudy (anonymous):

You are given two separate points: (x,14) and (18,y) You need to find out what the x and the y are though. To do this you have a handy relation between x and y. Namely that: y = 0.5x + 12 So if you are given the y value (14 for the first point) you can plug in that value and find what the x must be. Then you do the same thing for the second point, but you are given the x value and must find the y.

OpenStudy (anonymous):

omg. 2 / 1/2 = 4 doesn't it. I'm so useless without my calculator

OpenStudy (anonymous):

now do I do y = 0.5(18) + 12? so y = 9+12? so y = 21 OH and the book says yes. Thank you polpak.

OpenStudy (anonymous):

\[\frac{2}{\frac{1}{2}} = \frac{2}{\frac{1}{2}} \cdot 1= \frac{2}{\frac{1}{2}} \cdot \frac{2}{2}= \frac{4}{1} = 4\]

OpenStudy (anonymous):

Yep. Good job.

OpenStudy (anonymous):

your work here is done. Go now. Help the world.. you beautiful guru.